OpenStudy (anonymous):

Evaluate F(x) = 1/3x for x=4

OpenStudy (anonymous):

If x = 4 you would plug it into the equation to get f(x)

OpenStudy (anonymous):

i know thats what i did and nun of the answers are there to the question

OpenStudy (anonymous):

so i dont know if im plugging it in wrong? or something like i dont know?

OpenStudy (anonymous):

f(x) = 1/3 times 4. F(4) = 4/3

OpenStudy (anonymous):

Thats not and opinion tho its only A.-12 B.1 1/3 C.3/4 D.1/12

OpenStudy (anonymous):

so 3/4?

OpenStudy (anonymous):

4/3 = 1 1/3

OpenStudy (anonymous):

thank you !